# Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model

### v1

@@ 5 Discussion and conclusions
In this paper, we present predictions of the latest version of the gaea model, run on merger trees extracted from the p-millennium simulation. We compare model results against a compilation of observational measurements of galaxy clustering as a function of stellar mass and star formation rate, covering a wide redshift range from the local Universe up to the highest redshift when this measurement has been possible (i.e. z3.5). Our model shows a satisfactory agreement with the observed projected 2pCF, of comparable or superior quality with respect to independent semi-analytic codes (L-galaxies Guo et al. 2011; Henriques et al. 2017) and hydro-dynamical simulations (Springel et al., 2018; Artale et al., 2017). At z0, gaea results are close to those predicted from the TNG hydro-simulation suite (Springel et al., 2018): a good agreement with data is found at intermediate-to-low stellar mass bins and at the smallest scales, while the 2pCF in the highest mass bin has a smaller amplitude with respect to observational estimates. Similar results have been also found for predictions from the EAGLE simulation (Artale et al., 2017). Predictions from the L-galaxies model are able to better reproduce the amplitude in the highest-mass bin, but they either overpredict (Guo et al., 2011) or underpredict (Henriques et al., 2017) the clustering of intermediate-to-low mass galaxies. Overall, our model does not predict a significant evolution of the clustering strength with redshift, in good agreement with independent theoretical predictions (Artale et al., 2017) and observational measurements up to z1 (Marulli et al., 2013). At higher redshift, data seem to suggest a more significant evolution of the clustering amplitude than model predictions, although these are still in agreement with data within 1- errorbars and cosmic variance in the observed samples might play a relevant role. More data at z2 are definitely needed to clarify if the discrepancy between available measurements and model predictions is significant.
We estimate the clustering length corresponding to our mock samples, by considering the value corresponding to a 2pCF . gaea predicts values in the range 4 to 6 h-1 Mpc, mildly increasing with stellar mass. These values are in good agreement with observational constraints for intermediate-to-low mass galaxies (see e.g. Foucaud et al. 2010; Wake et al. 2011; Lin et al. 2012), but underestimate the expected values at M1011 h-1 M⊙ (typically larger than 10 h-1 Mpc). This is a reflection of the underestimate of the clustering amplitude in the most massive stellar mass bin that we discussed in Sec. 4. At fixed stellar mass, we find only a weak evolution of with redshift, in agreement with the numerical results from Artale et al. (2017).
If we split our model galaxies in star-forming and quiescent samples, we find that gaea correctly reproduces the increased clustering of red quiescent galaxies with respect to blue star forming systems. At M1011 h-1 M⊙, the agreement with data for the blue population is excellent, while red galaxies show a clustering excess at intermediate stellar masses. This shows that the modelling of physical mechanisms responsible for quenching in gaea (both internal and environmental) is able to reproduce the observed clustering strength and its variation with stellar mass. This is not a trivial success for our model, and it shows that our approach is not only able to recover the global properties of galaxy populations in a statistical sense, but also their spatial distribution in the LSS. For the most massive galaxies (larger than 1011 h-1 M⊙), both the clustering amplitude of the quiescent and star-forming populations are underestimated. This is consistent with the underestimate we find for the total population, and represents the main tension we found.
Most of the model galaxies contributing to the most massive bin are central galaxies. The underprediction of clustering amplitude for these galaxies highlights a mismatch between the predicted and expected mass of the host DMHs: gaea galaxies tend to live in DMHs slightly less massive than what suggested by the clustering analysis. On the other hand, at low-stellar masses, that are dominated by the satellite population, the good agreement between data and predictions for the 1-halo term (i.e. at scales 2-3 Mpc) suggests that the satellite distribution within halos is correctly reproduced. This implies that our treatment for the evolution of satellite galaxies as in Xie et al. (2020) is effective in recovering the main properties of this population, both in terms of stellar mass and star formation activity. Conversely, the excess of clustering signal at low-masses for the 2-halo term ((i.e. at scales 2-3 Mpc) suggest that gaea is predicting a larger than expected number of satellites in massive haloes: this can be related to our treatment of the merging times for orphan galaxies (i.e. satellite galaxies whose substructure has fallen below the resolution limit of the simulation).
@@
In summary, the latest gaea version shows a remarkable level of agreement for model predictions run on mergers trees extracted from simulation of different resolution and (slightly) different CDM cosmological parameters. This result is driven by the interplay between the new treatment for AGN (and its feedback on the host galaxies) and the explicit partitioning of cold gas into its atomic and molecular components. Both physical mechanisms act as regulators of the cold gas in model galaxies. gaea reproduces well the observed dependence of galaxy clustering on stellar mass and star formation activity. Moreover, model predictions are also consistent with the clustering measurements corresponding to HI-selected galaxies. This is an important confirmation that our treatment of the mechanisms acting on the baryonic component of galaxies, and satellites in particular, is able to capture the relevant physical dependencies. gaea predicts a small redshift evolution of the amplitude of the 2pCF: this is consistent with available data up to z3.5, which we interpret as another indication that our model is able to correctly connect galaxies with their environment. At higher redshifts, our model tends to under-predict the observational measurements. However, cosmic variance may play an important role at these redshifts and larger observational samples are needed to draw firmer conclusions in this redshift regime.
@@ Acknowledgements.
An introduction to gaea , a list of our recent work, as well as datafile containing published model predictions, can be found at https://sites.google.com/inaf.it/gaea/home. We acknowledge the use of INAF-OATs computational resources within the framework of the CHIPP project (Taffoni et al., 2020) and the INAF PLEIADI program (http://www.pleiadi.inaf.it). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We thank Chen Li for sharing his observational estimates for the clustering signal in the SDSS and DEEP2 surveys. We acknowledge stimulating discussions with Andrea Biviano, Pierluigi Monaco and Emiliano Sefusatti. MH acknowledges funding from the Swiss National Science Foundation (SNSF) via a PRIMA grant PR00P2-193577 ‘From cosmic dawn to high noon: the role of BHs for young galaxies’.References
@@
- Wang et al. (2013) Wang, L., Weinmann, S. M., De Lucia, G., & Yang, X. 2013, MNRAS, 433, 515

### v2

@@ 5 Discussion and conclusions
In this paper, we present predictions of the latest version of the gaea model, run on merger trees extracted from the PMS simulation. We compare model results against a compilation of observational measurements of galaxy clustering as a function of stellar mass and star formation rate, covering a wide redshift range from the local Universe up to the highest redshift when this measurement has been possible (i.e. z3.5). Our model shows a satisfactory agreement with the observed projected 2pCF, of comparable quality with respect to independent semi-analytic codes (L-galaxies Guo et al. 2011; Henriques et al. 2017) and hydro-dynamical simulations (Springel et al., 2018; Artale et al., 2017). At z0, gaea results are close to those predicted from the TNG hydro-simulation suite (Springel et al., 2018): a quite good agreement with data is found in the 10 ¡ M⋆ ¡ 10 stellar mass range and at the smallest scales, while the predicted 2pCT slightly under-predicts observational measurements for the most massive galaxies. Similar results have been also found for the EAGLE simulation (Artale et al., 2017). Predictions from the L-galaxies model are closer to the observed amplitude in the highest-mass bin, but they either overpredict (Guo et al., 2011) or underpredict (Henriques et al., 2017) the clustering of intermediate-to-low mass galaxies (i.e. 10 M 10) At the lowest mass bin, our model predictions over-predict the clustering signal, similarly to what found by Guo et al. (2011). Overall, our model does not predict a significant evolution of the clustering strength with redshift, in good agreement with independent theoretical predictions (Artale et al., 2017) and observational measurements up to z1 (Marulli et al., 2013). At higher redshift, data seem to suggest a more significant evolution of the clustering amplitude than model predictions, resulting in an overprediction with respect to the observed normalization of the 2pCF. Considering the observational uncertainties, model predictions are still consistent with observational data, and more data at z2 are needed to clarify if the highlighted discrepancy is significant.
Fig. 7 shows the dependence of the clustering lenght on galaxy stellar mass and its redshift evolution as predicted by our runs, and compares it to observational measurements. We compute from the intrinsic three-dimensional 2pCF (), directly computed from our model outputs, defined as the scale corresponding to . gaea predicts a significant increase of as a function of stellar mass. For M10 the predicted values are in reasonable agreement with observational constraints (see e.g. Foucaud et al. 2010; Wake et al. 2011; Lin et al. 2012; Marulli et al. 2013), with a clear excess of the predicted value of r0 for galaxy stellar masses below 10. At fixed stellar mass, we find only a weak evolution of at , in agreement with the numerical results from Artale et al. (2017) and Springel et al. (2018). A significant discrepancy between model predictions and data is found at z2, with gaea predicting a factor of 2 increase of at fixed stellar mass, and data exhibiting a slower trend as a function of galaxy stellar mass evolution.
If we split our model galaxies in star-forming and quiescent samples, we find that gaea correctly reproduces the increased clustering of red quiescent galaxies with respect to blue star forming systems. Overall, the red population shows a better agreement with data over the entire galaxy mass range considered, while blue galaxies exhibit both an excess of the clustering signal for low stellar masses (M109 M⊙) and an underprediction of the 2pCF at large masses (M1011 M⊙). Both discrepancies are consistent with the tensions we find between model prediction and data for the global population. In general, gaea results align well with similar comparisons shown in Guo et al. (2011) and Springel et al. (2018) for an independent semi-analytic and hydro-dynamical model of galaxy formation. These models tend to better reproduce the properties of the star-forming population, while overpredicting the clustering of the red population at intermediate stellar mass (in particular, for M 109-1010 M⊙).
Finally, we also compare our model predictions with clustering measurement for HI-selected galaxies in the ALFA survey and we find a good agreement over the entire mass range considered. Overall, the results discussed in this paper demonstrate that the proposed modelling of physical mechanisms responsible for gas consumption, recycling and quenching in gaea (both internal and environmental) are able to capture the main trends found for the observed clustering strength and its variation with stellar mass, SF activity and HI content. We stress that our model has not been calibrated to reproduce galaxy clustering, that thus represents a genuine model prediction: this is not a trivial success, and it shows that our approach is not only able to recover the global properties of galaxy populations in a statistical sense, but also their spatial distribution in the LSS. The general good agreement between model predictions and data for the 1-halo term (i.e. at scales 2-3 Mpc) for stellar masses down to M 109 M⊙, suggests that the satellite distribution within halos is correctly reproduced, at least down to this stellar mass limit. This implies that our treatment for the evolution of satellite galaxies as in Xie et al. (2020) is generally effective in recovering the main properties of this population, both in terms of stellar mass and star formation activity. The excess of clustering signal found for lower stellar masses can be understood as a combination of different effects. On the one hand, satellite galaxies residing in low-mass haloes are over-massive with respect to what is found in subhalo abundance matching models constrained to reproduce both the 2pCF and the GSMF (Wang et al., 2013a). The clustering signal excess for the 2-halo term ((i.e. at scales 2-3 Mpc) suggests that gaea is predicting a larger than expected number of satellites in massive haloes and this can be related to our treatment of the merging times for orphan galaxies.
@@
In summary, the latest gaea version shows a remarkable level of agreement for model predictions run on mergers trees extracted from simulations of different resolution and (slightly) different CDM cosmological parameters. This result is driven by the interplay between the new treatment for AGN (and its feedback on the host galaxies) and the explicit partitioning of cold gas into its atomic and molecular components. Both physical mechanisms act as regulators of the cold gas in model galaxies. gaea reasonably reproduces the observed dependence of galaxy clustering on stellar mass and star formation activity. Moreover, model predictions are also broadly consistent with the clustering measurements corresponding to HI-selected galaxies. This is an important confirmation that our treatment of the mechanisms acting on the baryonic component of galaxies, and satellites in particular, is able to capture the relevant physical dependencies. gaea predicts a small redshift evolution of the amplitude of the 2pCF: this is consistent with available data up to z3.5, which we interpret as another indication that our model is able to correctly connect galaxies with their environment. At higher redshifts, our model tends to under-predict the observational measurements. However, larger observational samples are needed to draw firmer conclusions in this redshift regime.
@@ Acknowledgements.
An introduction to gaea , a list of our recent work, as well as datafile containing published model predictions, can be found at https://sites.google.com/inaf.it/gaea/home. We acknowledge the use of INAF-OATs computational resources within the framework of the CHIPP project (Taffoni et al., 2020) and the INAF PLEIADI program (http://www.pleiadi.inaf.it). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We thank the anonymous referee for a careful reading of the manuscript that helped us improve the quality of the presentation of our results. We thank Chen Li for sharing his observational estimates for the clustering signal in the SDSS and DEEP2 surveys. We acknowledge stimulating discussions with Andrea Biviano, Alessandra Fumagalli, Federico Marulli, Pierluigi Monaco and Emiliano Sefusatti. MH acknowledges funding from the Swiss National Science Foundation (SNSF) via a PRIMA grant PR00P2-193577 ‘From cosmic dawn to high noon: the role of BHs for young galaxies’.References
@@
- De Lucia et al. (2010) De Lucia, G., Boylan-Kolchin, M., Benson, A. J., Fontanot, F., & Monaco, P. 2010, MNRAS, 406, 1533
@@
- Marulli et al. (2016) Marulli, F., Veropalumbo, A., & Moresco, M. 2016, Astronomy and Computing, 14, 35
@@
- Wang et al. (2013a) Wang, L., De Lucia, G., & Weinmann, S. M. 2013a, MNRAS, 431, 600
- Wang et al. (2013b) Wang, L., Weinmann, S. M., De Lucia, G., & Yang, X. 2013b, MNRAS, 433, 515